In an experiment to measure the effectiveness of a type of gas angles a fraction was found to be 31° in the angle of incidence is 55° calculator refractive index of the class


1
Expert's answer
2022-04-15T16:40:05-0400

"n=\frac{\sin\beta}{\sin\alpha}=1.6."